OpenStudy (anonymous):

HELP?!?!?!?!?!?!? Solve by using the perfect squares method. x^2 + 4x + 4 = 0 A. 2 B. 4 C. –4 D. –2

OpenStudy (anonymous):

can you help me?

OpenStudy (anonymous):

@AkashdeepDeb

OpenStudy (anonymous):

@paki

OpenStudy (akashdeepdeb):

x^2 + 4x + 4 = 0 Well to find the values of x when that whole expression equals zero. You need to know some rules. \[a^2 + b^2 + 2ab = (a+b)(a+b) = (a+b)^2\]\[a^2+b^2-2ab = (a-b)(a-b) = (a-b)^2\] If you notice, both the equations there ^ are perfect squares, which IS what we want! :D So x^2 + 4x + 4 = 0 is written in the first form. Thus,\[x^2 + 2(2)(x) + 2^2 = 0 ~~~~~~~~~~(Written~~in~~the~~proper~~format)\] \[(x+2)(x+2) = (x+2)^2 = 0\] Thus, for (x+2)^2 to be equal to zero. x has to be -2,-2 [which are the two solutions or roots]

OpenStudy (anonymous):

omg thank you!!! but i thought the formula was a^2+bx+c=0?

OpenStudy (anonymous):

sorry ax^2+bx+c=0

OpenStudy (anonymous):

@amistre64

OpenStudy (akashdeepdeb):

Yes, that IS the formula of a quadratic equation! And this is also a quadratic equation. The rules that I wrote up there were just identities, Say, if I have x^2 - 6x + 9 = 0 You can use one of those rules [the 2nd rule] and write it like: x^2 - 2(3)(x) + (3)^2 = (x-3)^2 = 0 Thus x = 3,3

OpenStudy (anonymous):

thank you!!!

OpenStudy (akashdeepdeb):

:)